# Accurate Checkerboard Corner Detection under Defoucs

### v1

@@ Abstract
Camera calibration is a critical process in 3D vision, impacting applications in autonomous driving, robotics, architecture, and so on. This paper focuses on enhancing feature extraction for checkerboard corner detection, a key step in calibration. We analyze existing methods, highlighting their limitations and propose a novel sub-pixel refinement approach based on symmetry, which significantly improves accuracy for visible light cameras. Unlike prior symmetry based method that assume a continuous physical pattern, our approach accounts for abrupt changes in visible light camera images and defocus effects. We introduce a simplified objective function that reduces computation time and mitigates overfitting risks. Furthermore, we derive an explicit expression for the pixel value of a blurred edge, providing insights into the relationship between pixel value and center intensity. Our method demonstrates superior performance, achieving substantial accuracy improvements over existing techniques, particularly in the context of visible light camera calibration. Our code is available from https://github.com/spdfghi/Accurate-Checkerboard-Corner-Detection-under-Defoucs.git.
@@ 1 Introduction
In intrinsic calibration, the projection camera model with non-linear distortion terms commonly used today was proposed by Brown [6, 8], while Zhang established an effective method for calibration based on 2D plane point mapping [33], which is widely applied. Additionally, there are calibration methods based on line distortion correction [5] and general calibration methods [18, 25] that fully consider point and line mapping. Feature extraction is the first step of calibration, which ultimately determines the accuracy and is the focus of this paper. The two main types of calibration targets are checkerboards and circles; checkerboards have the advantage of higher robustness to distortion, while circles offer higher accuracy [15, 14, 4]. The main issue with the circles pattern is that the shape of the circle under distortion can hardly be calculated. Recently, Song et al. [28] proposed an impressive analytical method to compute the centroid of a distorted circle and outperforms checkerboard-based calibration. But this method still has limitations, as extracting centroid in the image requires a larger pattern area than traditional circle fitting method, thus restricting the position of the circle’s center from the image edge, where distortion is most significant.
@@
To improve the detection accuracy of checkerboard corner points, a series of methods have been studied. Currently, two mainstream approaches are notable. The first is assuming vector from a corner to its adjacent area pixel is orthogonal to pixel’s gradient direction, such as implemented in OpenCV [10], which achieve good sub-pixel accuracy; however, its performance is subpar under some conditions, as tested by Schöps et al. [27] and Song et al. [28] with infrared cameras. Datta et al. discovered that the assumption of OpenCV does not hold in non fronto-parallel input images [4]. Therefore they iteratively estimate the homography matrix, and then extract point after a transform to canonical fronto-parallel image, which result in a significantly improvement. The second approach involves quadratic surface fitting [20], commonly used in MATLAB. Fitting methods not only apply to regular checkerboards but have also been extended to curved surface[31]. Ha et al. [12] expanded the fitting method to patterns with three intersecting lines, and got better reprojection error than checkerboard which has two intersection lines. Schöps et al. [27] further developed Ha’s pattern by using a star-shaped calibration board with eight intersecting lines, they proposed a novel feature extraction algorithm for the pattern [26, 27], this method takes cost by pixel differences in projections of two world points that are symmetric about pattern center. With more gradient information, they got significant improvement in accuracy. They tested their method on the expanded checkerboard pattern using both RGB and infrared cameras, achieving reprojection errors that was reduced to and , respectively, compared to the OpenCV method, which is impressive.
@@
However, we are curious why their method shows such a significant difference between the RGB and infrared cameras, and is it possible that their method is not the most effective for the RGB cameras. We carefully considered their assumptions. First, their optimization objective function implicitly assumes that the image is captured by a pinhole camera of a continuous physical pattern, and that there should not be significant distortion. Then, they assume that the image vary smoothly among neighbor pixels. Regarding the former, we believe that the assumption of a continuous physical pattern does not always hold. For the infrared cameras, since infrared radiation is based on the spontaneous emission from the thermal motion of molecules, and it is almost impossible for the surface temperature of an object to change abruptly, the assumption of a continuous spatial pattern is reasonable. However, for the RGB cameras, which captures light reflected by objects and is related to the object’s color, abrupt changes can occur within the range of a single pixel[29]. For blurry images, defocus becomes the dominant factor. Thus, we proposed a novel objection function by consider this kind of blur, experiment of visible light camera shows agree with our assumption. Moreover, our objective function is simpler with fewer parameters, which means reduced computation time and a lower risk of overfitting. We also do a math on their smooth assumption by study a blur line model, which is the convolution of 2D step function with a Gaussian kernel. The closest model to the one we used is Ref. [1], where they addressed constructing feature detector. We provided explicit expression for pixel value the first time and used it to study the variation of differences against intensity of pixel center.
@@
We propose a novel sub-pixel refinement method for checkerboard corner detection based on symmetry and provide a detailed analysis of its physical assumptions. This method shows significant accuracy improvements in visible light camera compared with existing symmetry based method and mitigates overfitting risks.
For the first time, we derive an explicit expression for the pixel value (as a region integral) of an blur edge based on the convolution model of a 2D step function with a Gaussian kernel. Using this analytical result, we analyze the difference between the pixel value and the intensity at the pixel center.
@@ 2 Related Works
Camera Calibration under defocus Duda et al. [7] proposed a method based on a localized Radon transform approximated by box filters, compared to the OpenCV method, it is more robust to noise and blur, achieving significant accuracy improvements in outdoor environments. Ha et al. [11] introduced an active targets approach for calibration under defocus, they separated the checkerboard into overlapping horizontal and vertical stripes, and then applied a novel deconvolution method to these stripes. Further, they do a refraction correction of the glass screen. This method was evaluated on real data, and shows accurate camera calibration result under severe defocus.
@@
Line Detection under Defocus. The classic Canny algorithm obtains single-pixel edges through non-maximum suppression, and Devernay et al. further refines edges to sub-pixel accuracy using quadratic fitting near maxima[30]. While these well-known methods assume edge is maxima of smooth function, Ref. [29] first considers the physical scenario of a straight edge partially falling within a pixel. By analytically integrating the area segmented by the line within one pixel, it shows the discontinuity of discrete pixels will generate bias with traditional gradient computation. But they didn’t provide an analytical result considering defoucs. The calibration pattern used by Ha et al. [11] can actually be considered as lines under defocus, where they proposed an deconvolution method but also without using analytical formula. The analytical expression for line defocus was first introduced in Ref. [22] and later used in Ref. [35] for single-image defocus estimation. Bakery et al. [1] considered the pixel value rather than point value expression, which was not addressed in Ref. [22]. The line defocus model used in this paper is similar to Bakery et al., the key difference is that their paper focused on using numerical methods to construct a feature detector. They only provided an integral form while we provide explicit results.
@@ 3 Method
3.1 Camera model
We briefly review the pinhole camera and distortion model, which will be used in our experiment part. For a point in homogeneous representation of camera coordinate, the distorted point is represented as [8]
and
where
are radial distortion coefficients and are tangential distortion coefficients.
For modern, well-manufactured cameras, the shear deformation component can be ignored, and the camera model takes the following form [13]
the image coordinates can be obtained as
3.2 Line defocus model
Before introducing our algorithm, we would like to discuss pixel continuity and the effect of image blur on interpolation. In the absence of blur, traditional subpixel techniques tend to produce significant deviations near image boundaries, which is the so called partial area effect[29]. A simple example involves two horizontally adjacent pixels, and , with a vertical boundary dividing pixel into two equal area, where the intensities of the two parts are and , respectively, while the intensity of all positions in pixel is . Using linear interpolation yields an intensity of at the location between the two pixels, which differs significantly from the true intensity . Intuitively, Gaussian blur can reduce this discrepancy, and we will provide a quantitative discussion of this effect in this section.
@@ 3.2 Line defocus model
where W and B denote constant intensity of white and black region, and d represents the distance from point to the line passing through the feature center . Let denotes the angle from image x-axis to the line, one can compute through a simple cross product
Note that the two half-planes divided by the line are not identical, so is a signed quantity.
@@
The real blur image can be calculated as a convolution of noise-free pinhole image with the kernel. Due to the symmetry of Gaussian function, we align the y-axis with direction of the line (Fig. 2(b)-(d)), which simplifies the convolution to a straightforward integration
@@
Fig. 3 shows the difference between Eq. (13) and Eq. (9), where letter use the center position of pixel. We randomly select the tilt angle of the line and randomly vary the position of within the range of one pixel. It can be observed that, under typical image parameters, the maximum difference between the pixel center intensity and the pixel value becomes negligible when approaches one pixel. This condition is easily satisfied during calibration.
3.3 Refinement via symmetry
For the formulas provided in the previous section, a direct approach is to perform nonlinear optimization on the pixel values. Considering that the checkerboard corners are the intersection points of two lines, for the effects of aliasing in checkerboard corner’s center, though the pattern using a display device proposed by Ha et al. [11] can be used to mitigate it. However, using a display is not the optimal choice in all calibration scenarios, and the refraction correction of display screen may introduce extra errors.
@@ 3.3 Refinement via symmetry
where is the position in world coordinate centered at every pattern center, and is the homography that projects the world coordinate into image pixels. By using the Levenberg-Marquardt(LM) algorithm, they estimates the homography matrix for every feature, the result points was derived from the homography matrix. This method makes amazing result. For eight intersecting line pattern, they reduce the medium reprojection errors to less of OpenCV method with infrared camera.
However, we believe this method has two limitations: Firstly, Eq. (16) implicitly assumes that the image is captured by a pinhole camera of a continuous physical pattern, this holds for infrared camera since the surface temperature of an object can hardly change abruptly. But for visible light cameras, which captures light reflected by objects, abrupt changes can occur within the range of a single pixel[29]. In calibration, the blur in image is mainly caused by defoucs, see Fig. 4. Secondly, distortion may also leads to a poor initialization of LM algorithm. Moreover, although straight lines can still be considered straight in a small neighborhood despite distortion, which is why the chessboard method is effective. But when number of intersection lines large than two, the mapping of these lines may not be described by a homography under distortion.
Now let’s do a more physical modification of Eq. (16) for visible light camera. Within the pixel ranges used for calibration, the Gaussian of kernel function can be considered a constant. This assumption works well even with large defoucs up to several tens of pixels [12], though in that case we should consider the environment light change. We cannot use a single homography matrix to describe this mathematical relationship. A homography matrix has eight degrees of freedom and can be determined by the correspondence of four pairs of points. Considering two intersecting lines in a checkerboard pattern, a homography matrix can be computed by selecting a bounding box around any one of the blur lines. However, the pixels near the other blur line are not necessarily determined by this matrix. In fact, the other line can be at any arbitrary angle. But from Eq. (9), we can see that as long as two points are on the same side of the half-plane and equidistant from the boundary line, they have the same value. Furthermore, when another line exists, i.e., in the standard checkerboard pattern, if two points are symmetric with respect to the feature point , then they have the same value. Therefore, we can formulate the following optimization objective
@@
4 Experiment
4.1 Setup
Many previous studies have achieved excellent results on synthetic data but performed a little bit poorly on real data; therefore, we directly use real data for evaluation. To date, there is no widely recognized calibration test dataset that includes the ground truth of camera parameters, so we make our own like other researchers. We primarily evaluate our results on two datasets: one captured by the infrared camera of an Intel RealSense D435 and visible light camera HIKVISION MV-SC2005AM. The latter has more noise and distortion compared to the former. The reason we chose latter is that RGB cameras may exhibit purple fringing, which could affect our assumptions. In this paper, we focus more on the impact of defocus on calibration, so we selected a camera that captures grayscale images. Besides, we want to know how distortion will affects the algorithms we compare. We evaluate accuracy using the widely used reprojection error, which projects world points into image coordinates using estimated camera parameters and then calculate the difference of extracted feature coordinates and projected positions.
We compared the performance of three methods: OpenCV, Ref. [27] and ours. Here, we are not concerned with the success rate of detection, although that is another topic worth discussing [23, 9]. We focus solely on the accuracy of sub-pixel refinement, thus using initial feature values obtained from OpenCV. Due to variations among different calibration optimization algorithms, we consistently use OpenCV’s implementation for all feature refinement methods, which initializing with Zhang’s method [33] and employing a third-order radial and second-order tangential distortion model[8] as Eqs. 1 and 2, followed by LM method up to 30 iterations.
4.2 Results
We have noticed that few studies discuss the impact of the optimization window on accuracy in refinement algorithms. Some, like Refs. [7] and [27], mention using a fixed window for their experimental data. However, we observe that, especially in cases with large distortion and perspective effects, the impact of the optimization window size on accuracy can exceed the differences between algorithms. In Fig. 5, we show the variation of reprojection errors with optimization window size for different algorithms. Fig. 5(a) presents the results for an infrared camera, as shown in Tab. 2, the distortion of this camera is relatively small. Therefore, all algorithms achieve the best reprojection error over a large window size range. In the range tested in our experiments, methods of this paper and Ref. [27] significantly outperform OpenCV’s method. The reprojection error for OpenCV stabilizes as the window size increases, while both methods of this paper and Ref. [27] show a slight decrease from to , followed by a slow rise, with Ref. [27]’s method rising more slowly. As seen in Tab. 2, for the optimal reprojection error, our method is similar to Ref. [27]. Compared to OpenCV, our method reduces the mean and median reprojection errors by and , respectively.
Fig. 5(b) shows the results for visible light camera. Compared to the infrared camera we used, it has a lower resolution and a wider field of view. As seen in Tab. 2, it has larger camera distortion. In this case, the accuracy of both the OpenCV and Ref. [27]’s methods fluctuates significantly with the change in window size. In particular, Ref. [27]’s method shows an overall decrease, but with a fluctuation around size . We believe this is mainly due to the inappropriate physical assumptions of this method in this situation, as discussed earlier. Additionally, we observed some sample points fell outside the image boundaries during the algorithm’s execution, which is caused by difficulties in initializing the homography matrix due to distortion and noise, while the LM algorithm is sensitive to initial values. We discarded these points. However, the same algorithm did not encounter this issue in any of the windows for the infrared camera. For OpenCV, its overall error decreases first and then increases, achieving good reprojection error only in the narrow window size range of . When the window size increases to , Ref. [27]’s method reduces the error by almost compared to OpenCV. For the optimal error, we significantly outperform both methods, reducing the mean and median reprojection errors by and compared to Ref, and by and compared to OpenCV.
| Reprojection error | Camera matrix | Distortion() | |||||||||
| Methods | Average | Medium | |||||||||
| Infrared camera | |||||||||||
| OpenCV | 0.0842 | 0.0635 | 640.55 | 642.15 | 643.42 | 400.23 | 4.82 | -10.00 | 5.85 | 0.07 | 0.35 |
| Proposed | 0.0576 | 0.0430 | 640.75 | 642.34 | 643.54 | 400.31 | 4.30 | -7.55 | 4.15 | 0.14 | 0.42 |
| Ref. [27] | 0.0561 | 0.0420 | 640.80 | 642.41 | 643.52 | 400.33 | 4.34 | -7.60 | 4.19 | 0.16 | 0.40 |
| Visible light camera | |||||||||||
| OpenCV | 0.0848 | 0.0627 | 471.07 | 470.93 | 386.14 | 296.41 | 27.53 | -74.05 | 16.00 | -0.19 | -0.12 |
| Proposed | 0.0636 | 0.0470 | 471.30 | 471.17 | 386.24 | 296.43 | 24.88 | -68.17 | 11.76 | -0.19 | -0.09 |
| Ref. [27] | 0.0870 | 0.0577 | 471.29 | 471.16 | 386.04 | 296.47 | 27.71 | -76.89 | 16.34 | -0.17 | -0.29 |
Tab. 2 presents the estimated parameters, distortion coefficients are from Eqs. 1 and 2. For the infrared camera, our results are similar to those of the Ref. [27]. For the visible light camera, the camera matrix parameters estimated by the Ref. [27] are similar to ours, while the distortion parameters align more closely with those of OpenCV. Fig. 6 shows the distribution of the reprojection error. In the visible light camera case, the reference method produces large errors in several views. In the blue-colored view, outliers are observed in our method, which corresponds to the viewpoint with the highest reprojection error using the OpenCV method. The primary cause is that, in this view, some feature points are significantly farther from the camera’s z-axis, leading to higher noise levels. While this issue can be solved through such as manual correction by MATLAB toolbox, we decided to keep the original outcomes of algorithm.

### v2

# Feature Extraction Reimagined: Achieving Superior Accuracy in Camera Calibration

@@ Abstract
Camera calibration is crucial for 3D vision applications. This paper focuses on improving the accuracy of feature extraction, which is a key step in calibration. We address the aliasing problem of star-shaped pattern by introducing a novel dynamic calibration target that synthesizes multiple checkerboard patterns of different angle around pattern center, which significantly improves feature refinement accuracy. Additionally, we propose a novel cost function of feature refinement that accounts for defocus effect, offering a more physically realistic model compared to existing symmetry based method, experiment on a large dataset demonstrate significant improvements in calibration accuracy with reduced computation time. Our code is available from https://github.com/spdfghi/Feature-Extraction-Reimagined-Achieving-Superior-Accuracy-in-Camera-Calibration.git.
@@ 1 Introduction
Camera intrinsic calibration is mainly divided into three steps: feature extraction, parameter initialization, and nonlinear optimization. Camera models can be classified into two categories: physics-based parametric models and purely mathematical generic models. In physical models, geometric calibration assumes that the camera generally follows the pinhole camera projection model, while also considering the nonlinear effect of the lens on the light rays, leading to what is known as distortion. The two main types of distortion models are the Brown-Conrady model [5, 7] and the Thin-Prism Fisheye model [31]. Generally, during model initialization, distortion is ignored, and an initial estimate of the camera matrix is obtained. Zhang et al. [32] proposed a initialization method based on the algebraic relationships of homography matrices between control points in multiple 2D planes, which is flexible and commonly used. After initialization, the nonlinear optimization step is used to calculate distortion parameters and refine the camera matrix elements. The widely adopted approach is to use sum of reprojection error norm as the cost function, there are also other kind of cost functions such as the one base on line straightness metric [4]. The generic camera model only considers the correspondence between pixel points and 3D observation lines[9, 16, 23, 25], Ref. [9] also consider parameter of defocus. The initialization of generic model is thoroughly considered in Ref. [23], Schöps et al. [25] extended their method, achieving more accurate feature extraction and provided full bundle adjustment (BA) optimization.
Parameter initialization and nonlinear optimization are purely mathematical problems that have been extensively explored, while feature extraction is still an open task. As the first step of calibration, feature extraction and refinement ultimately determines the accuracy in most cases and is the focus of this paper. The two main calibration pattern are the circle and checkerboard. The former generally provides higher accuracy but is limited by distortion, whereas the latter does not have this problem [13, 12, 3]. The main issue with the circle pattern is that the shape of circle under distortion can hardly be calculated. Recently, Song et al. [26] proposed an impressive analytical method to solve this problem, which computes the centroid rather than fitting the shape of a distorted circle. But this method still has limitations. For accurate centroid calculation, the circle must be sufficiently large, which restricts the ability to estimate for distortion near the image boundaries.
@@
To improve the detection accuracy of checkerboard pattern, a series of methods have been studied. Currently, two mainstream approaches are notable. The first is assuming vector from a corner to its adjacent area pixel is orthogonal to pixel’s gradient direction [8], such as implemented in OpenCV, which achieve good sub-pixel accuracy; however, its performance is subpar under some conditions, as tested by Song et al. [26] with infrared cameras. Datta et al. discovered that the assumption of OpenCV does not hold in non fronto-parallel input images [3]. Therefore they iteratively estimate the homography matrix and distortion, then extract point after a transform to canonical fronto-parallel image, which result in a significantly improvement. The second approach involves quadratic surface fitting [18], commonly used in MATLAB. Fitting methods not only apply to regular checkerboards but have also been extended to curved surface[29]. Ha et al. [11] expanded the fitting method to Deltille pattern which has three intersecting lines(Fig. 1), where they got better reprojection error than checkerboard. Schöps et al. [25] further developed Ha’s pattern by using a star-shaped calibration board with eight intersecting lines, they also proposed a novel feature refinement algorithm [24, 25], which takes cost by pixel differences in projections of symmetric world points. With more gradient information, they got significant improvement in accuracy.
However, their experiments show that best segment number range of star-shaped pattern is 12 to 20. Exceeding this number actually leads to an increase in reprojection error. They attribute this to the fact that monitor and camera cannot resolve more intersecting line patterns, i.e., the problem of aliasing caused by defocus and pixel discretization. To solve this problem, we employed a novel calibration pattern that consists of a video synthesized from a series of checkerboard patterns rotating around center of the pattern. By capturing these patterns at different time and then combining the captured images into a multi-channel image, we can separate areas that would otherwise experience aliasing into distinct layers, thereby solved this problem. Our experiments demonstrated the accuracy advantage against the optimal star-shaped pattern under same segmentation, with further improvements in accuracy as the segmentation increases.
@@
In addition, we carefully considered their cost function for feature refinement. First, it implicitly assumes that the image is captured by a pinhole camera of a continuous physical pattern, and that there should not be significant distortion. Then, they assume that the image vary smoothly among neighbor pixels. Regarding the former, we believe that the assumption of a continuous physical pattern does not always hold. Although infrared radiation is generated by molecular thermal motion, which may lead to blurred physical patterns in infrared-based cameras, this is not the case for commonly used visible light cameras, which captures light reflected by objects and is related to the object’s color. Abrupt changes can occur within the range of a single pixel [27], and defocus is the dominant factor for blur. Thus, we proposed a more physical cost function by considering of defocus. We conducted experiments on the calibration of a generic camera model using a dataset of over one thousand images, achieving significantly lower reprojection errors. Moreover, our cost function is simpler with fewer parameters, which means reduced computation time and a lower risk of overfitting. We also do a math on their smooth assumption by study a blur line model, which is the convolution of 2D step function with a Gaussian kernel. We provided explicit expression for pixel value and used it to study the variation of differences against intensity of pixel center.
@@
We propose a novel camera calibration system that addresses the aliasing issue in patterns based on multiple intersecting lines. This system provides superior calibration accuracy, while its shape remains unaffected by distortion, similar to the basic checkerboard pattern.
We propose a symmetry based cost function for feature refinement of checkerboard-like corner. This method shows significant accuracy and speed improvements compared with existing symmetry based approach.
@@ 2 Related Works
Camera Calibration under defocus Duda et al. [6] proposed a method based on localized Radon transform approximated by box filters, compared to the OpenCV method, it is more robust to noise and blur, achieving significant accuracy improvements in outdoor environments. Ha et al. [10] introduced an active targets approach for calibration under defocus, they separated the checkerboard into overlapping horizontal and vertical stripes, and then applied a novel deconvolution method to these stripes. Further, they do a refraction correction of the glass screen. This method was evaluated on real data, and shows accurate camera calibration result under severe defocus.
Their pattern shares some design similarities with ours, as both approaches decompose existing patterns into multiple simpler ones and then perform calibration dynamically. However, there are three notable differences: Firstly, their pattern is designed primarily to simplify deconvolution operations rather than address aliasing; they use only a decomposition of a checkerboard pattern, whereas our pattern, similar to star-shaped calibration board, provides more gradients. Secondly, they model defocus with a Gaussian function followed by deconvolution for feature refinement, while we leverage symmetry to implicitly account for defocus, making it more adaptable to various camera types. Lastly, our pattern does not rely on a display screen, by changing the pattern shape from square to circular, it can be manufactured mechanically, which is more flexible.
@@
Line Detection under Defocus. The classic Canny algorithm obtains single-pixel edges through non-maximum suppression, and Devernay et al. further refines edges to sub-pixel accuracy using quadratic fitting near maxima[28]. While these well-known methods assume edge is maxima of smooth function, Ref. [27] first considers the physical scenario of a straight edge partially falling within a pixel. By analytically integrating the area segmented by the line within one pixel, it shows the discontinuity of discrete pixels will generate bias with traditional gradient computation. But they didn’t provide an analytical result considering defoucs. The calibration pattern used by Ha et al. [10] can actually be considered as lines under defocus, where they proposed an deconvolution method but also without using analytical formula. The analytical expression for line defocus was first introduced in Ref. [21] and used in Ref. [34] for single-image defocus estimation. Bakery et al. [1] considered the pixel value rather than point value expression, which was not addressed in Ref. [21]. The line defocus model used in this paper is similar to Bakery et al., the difference is that their paper focused on using numerical methods to construct a feature detector. They only provided an integral form while we provide explicit results.
@@ 3 Method
3.1 Pattern design
Clearly, the aliasing of the star-shaped pattern in the captured images is caused by a combination of physical material properties, defocus, and pixel discretization. The first two factors are inevitable and occur in all kinds of patterns. Our design primarily addresses the latter, where multiple intersecting lines within a pixel range dominate. The key idea is that, unlike the star pattern board, which captures all intersecting line patterns in a single image, we instead extract gradient information from a series of checkerboard patterns rotated around the center of the pattern, see Fig. 2. These rotated-checkerboard images are then combined into a multi-channel image, i.e., we separate the information of spatially aliased pixels into different channels.
To obtain the checkerboard patterns rotated around the local centers, we cyclically display the calibration images from Fig. 3 on the screen. Then, from the video stream captured by the camera, we obtain stable frames based on the pixel differences between adjacent frames, and use the AprilTag to distinguish the rotation angles of different checkerboard patterns. These stable frames are combined into a multi-channel image, which is then used for feature optimization with the cost function from Sec. 3.3. We combine these images into a multi-channel format for two reasons: first, it integrates multiple calibration patterns, thereby mitigating system bias introduced by random sampling in our algorithm; second, it is cache-efficient, enhancing computational performance.
3.2 A Gaussian defocus model
Before introducing our cost function for feature refinement, we would like to discuss pixel continuity and the effect of image blur on interpolation. In the absence of blur, traditional subpixel techniques tend to produce significant deviations near image boundaries, which is the so called partial area effect[27]. A simple example involves two horizontally adjacent pixels, and , with a vertical boundary dividing pixel into two equal area, where the intensities of the two parts are and , respectively, while the intensity of all positions in pixel is . See Fig. 4. Using linear interpolation yields an intensity of at the location between the two pixels, which differs significantly from the true intensity . Intuitively, Gaussian blur can reduce this discrepancy, and we will provide a quantitative discussion of this effect in this section. It is worth noting that this issue also arises in the feature refinement problem where Datta et al. shows that OpenCV’s method doesn’t work well [3].
@@ 3.2 A Gaussian defocus model
where W and B denote constant intensity of white and black region, and d represents the distance from point to the line passing through the feature center . Let denotes the angle from image x-axis to the line, one can compute through a simple dot product
Where we build coordinate that y axis is along direction of the line and is the normal vector of x axis. Note that the two half-planes divided by the line are not identical, so is a signed quantity.
@@
The real blur image can be calculated as a convolution of noise-free pinhole image with the kernel. Due to the symmetry of Gaussian function, the convolution to a straightforward integration
@@
Fig. 6 shows the difference between Eq. (8) and Eq. (4), where letter use the center position of pixel. We randomly select the tilt angle of the line and randomly vary the position of within the range of one pixel. It can be observed that, under typical image parameters, the maximum difference between the pixel center intensity and the pixel value becomes negligible when approaches one pixel. This condition is easily satisfied during calibration, and is also necessary for our refinement algorithm.
3.3 Cost function in feature refinement
For the formulas provided in the previous section, a direct approach is to perform nonlinear optimization on the pixel values. Considering that the checkerboard corners are the intersection points of two lines, for the effects of aliasing in checkerboard corner’s center, though the pattern using a display device proposed by Ha et al. [10] can be used to mitigate it. However, not all cameras’ point spread functions can be approximated by a Gaussian function. In the generic camera model, elliptical Gaussians are also used as a defocus model [9], but in this case, it is difficult to obtain a simple analytical expression. Generally, a Gaussian approximation of the point spread function can only be applied in the cases of simple apertures and small defocus[19].
@@ 3.3 Cost function in feature refinement
where is the position in world coordinate centered at every pattern center, and is the homography that projects the world coordinate into image pixels. By using the Levenberg-Marquardt(LM) algorithm, they estimates the homography matrix for every feature, the result points was derived from the homography matrix. This method makes significant result better than OpenCV for star-shaped pattern.
However, we believe this method has two limitations: Firstly, Eq. (11) implicitly assumes that the image is captured by a pinhole camera of a continuous physical pattern, this may hold for infrared camera since the surface temperature of an object can hardly change abruptly. But for common used optical cameras, the blur in image is mainly caused by defoucs[9, 10, 34]. Secondly, distortion may also leads to a poor initialization of LM algorithm. Moreover, although straight lines can still be considered straight in a small neighborhood despite distortion, which is why the chessboard method is effective. But when number of intersection lines large than two, the mapping of these lines may not be described by a homography under distortion.
Now let’s do a more physical modification of Eq. (11). Firstly we assume that the Gaussian kernel is a suitable approximation for point spread function. Within the pixel ranges used for calibration, the Gaussian of kernel function can be considered a constant. This assumption works well even with large defoucs up to several tens of pixels [11], though in that case we should consider the environment light change. We cannot use a single homography matrix to describe this mathematical relationship. A homography matrix has eight degrees of freedom and can be determined by the correspondence of four pairs of points. Considering two intersecting lines in a checkerboard pattern, a homography matrix can be computed by selecting a bounding box around any one of the blur lines. However, the pixels near the other blur line are not necessarily determined by this matrix. In fact, the other line can be at any arbitrary angle. But from Eq. (4), we can see that as long as two points are on the same side of the half-plane and equidistant from the boundary line, they have the same value. Furthermore, when another line exists, i.e., in the standard checkerboard pattern, if two points are symmetric with respect to the feature point , then they have the same value. Therefore, we can formulate the following optimization objective
@@
An additional situation needs to be discussed: if the window size exceeds the physical size of the pattern(Fig. 8), our cost function may not yield good results, and the method of Schöps et al. would be more suitable. However, considering that we always set the optimization window small enough (around 20 pixels) in calibration for more accurate distortion calculations, and to better utilize the pattern information and avoid aliasing, we believe that in this case, the priority should be to adjust the physical size of the pattern.
4 Evaluation
4.1 Experiment Setup
Many previous studies have achieved excellent results on synthetic data but performed a little bit poorly on real data; therefore, we only use real data for evaluation. We evaluate our method on two datasets: one is the star-pattern dataset provided by Schöps et al., and the other consists of results captured using the Intel D435 color camera. Our data is obtained as follows: First, we convert the calibration pattern sequence from Fig. 3 into a PDF file, then use the built-in screen recording software on Windows to record the process of flipping through the PDF file, generating a calibration video. We then play this calibration video in a loop and capture the results using D435. From the captured video stream, we compare the average pixel difference between adjacent frames. If it is smaller than a threshold (set to 1 pixel), we consider the frame to be stable. From each stable frame sequence, we select the middle frame and determine the corresponding channel based on the AprilTag index. We are not concerned with the initialization of feature positions, as the initial positions of all features are determined by the star-shaped pattern. We use the generic camera calibration pipeline published by Schöps et al. on GitHub [25] for generic camera parameter calculation. This pipeline provides BA optimization for varies camera model including Brown-Conrady, Thin-Prism Fisheye and generic model.
We primarily evaluate accuracy using the direction of the reprojection error, the median reprojection error(MRE), and the KL-Divergence between scaled distribution of reprojection error and 2D normal distribution, which is used by Schöps et al. for metric of biasedness. Additionally, we compare the computational time difference between our cost function and that of Schöps et al. with no CUDA.
4.2 Evaluation of cost functions
To evaluate our cost function, we used the calibration data provided by Schöps et al. on GitHub(issue 16) [25], which includes 1379 images captured by a D435 infrared camera. Splitting the data into training and test sets is not a common practice for camera calibration evaluation, all data were used. We run the feature extraction and generic camera calibration pipeline provided by Schöps et al., with the parameter setting optimized for a window half-width of 10 pixels, and the cost function computed the image intensity difference.
| Method | MRE | KL | Average time |
| (pixels) | Divergence | (seconds) | |
| Schöps et al. | 2.3 | ||
| Proposed | 0.0326 | 0.49 | 0.41 |
From Tab. 2, our method reduces the MRS by compared to the method of Schöps et al., and the computation time is reduced to less than . Fig. 9 shows that both two methods achieve relatively uniform directions of reprojection error under the central generic model, without introducing strong systematic errors.
We do not expect significant differences between our method and that of Schöps et al.’s when the calibration pattern are mostly parallel to the camera plane. However, for datasets like those with many non-parallel images, and in cases where infrared illumination is not used, the D435I primarily captures visible light calibration patterns. In such situations, the advantages of our method can be anticipated.
4.3 Evaluation of calibration patterns
In the current implementation of pattern, the proposed scheme requires more time to capture each target during calibration compared to the static pattern. As a result, it is difficult to obtain thousands of calibration images by simply recording a video, making it challenging to calibrate a general camera model that requires dense data using manual methods. Therefore, we test using a simple Brown-Conrady model. We use 12 parameters, including 4 camera matrix elements, 6 radial distortion parameters, and 2 tangential distortion parameters. While this calibration method is implemented in OpenCV, considering the refraction correction issue mentioned by Ha et al.[10], we use the implementation of Schöps et al. [25] which including BA correction.
We evaluated three calibration patterns: one is the checkerboard pattern, or 4-segment star pattern; the second is 16-segment pattern proposed by Schöps et al., which showed the best calibration accuracy in their experiments; and the last is proposed pattern. Our pattern uses two different numbers of rotating checkerboard: one with 4, corresponding to the pattern proposed by Schöps et al., with the same black-and-white segments, and the other with 8.
| Pattern | Method | MRE | KL |
| (pixels) | Divergence | ||
| Star-4 | Proposed | 0.085 | 2.24 |
| Star-4 | OpenCV | 0.096 | 2.18 |
| Star-16 | 0.071 | 2.14 | |
| Mult-4 | Proposed | 0.066 | 2.16 |
| Mult-8 | 0.055 | 2.13 |
We summarized the experimental results in the Tab. 3, compared to the star-shaped calibration pattern with the same number of black-and-white segments, our method reduces the MRS by . Compared to using the checkerboard pattern and OpenCV feature refinement, our pattern with 8 rotating checkerboard reduces the reprojection error by . Our experiment used images, which is sufficient for parameterized camera model calibration. Additionally, we followed the approach of Schöps et al. by using a pattern with more control points than the standard chessboard. While typical experiments involve around points, each of our extracted feature points ranged from to . Such results are significant.